# CoSpaDi: Compressing LLMS via Calibration-Guided Sparse Dictionary Learning

### v2

@@ 4.2 Ablation Study
Figure 2: Dual-axis plot showing average accuracy (  solid lines, left axis) and perplexity (--- dashed lines, right axis, logarithmic scale with inverted direction) as functions of $\rho$ for Llama3.2-1B under three compression levels: 0.2, 0.3 and 0.4. Perplexity decreases upward due to axis inversion.
CR Bitwidth Avg. Acc. PPL
0.1686 bFP16 0.6198 1.94E+01
0.1843 bFP15 0.6195 1.95E+01
0.2001 bFP14 0.6176 1.97E+01
0.2726 bFP16 0.5408 4.34E+01
0.2864 bFP15 0.5394 4.37E+01
0.3002 bFP14 0.5373 4.46E+01
0.3765 bFP16 0.4096 1.80E+02
0.3883 bFP15 0.4086 1.82E+02
0.4002 bFP14 0.4069 1.84E+02
0.4804 bFP16 0.2846 8.23E+02
0.4902 bFP15 0.2838 8.15E+02
0.5001 bFP14 0.2837 8.10E+02
Table 1: Results on truncation of bfloat16 mantissa bits of coefficient matrix with reported average accuracy (Avg. Acc.) and Wiki Text word perplexity (PPL) for Llama3-8B resulting in different compression ratios (CR).
From the provided plots we can observe that depending on the compression ratio the optimal $ks$-ratio differs, thus, for simplicity in further experiments we select in all cases $\rho=2$.
Data-Free and Data-Aware Scenarios We argue that our proposed CoSpaDi allows for a more flexible representation compared to low-rank approximation, regardless of whether data-free or data-aware scenarios are considered. To prove this claim we performed an ablation study on Llama3-1B for different compression levels. We selected SVD for the data-free scenario and SVD-LLM for the data-aware case, while varying the compression ratio from 0.2 to 0.5. In Table 4.2 we report the results only for 0.2-0.5 compression. These results clearly indicate that in both data-free and data-aware settings the SDL-based methods outperform low-rank baselines by a wide margin, while CoSpaDi outperforms all methods.
@@
SVD x 0.5201 0.2566 0.0003 0.2471 0.2210 0.2000 0.2144 0.2537 0.2391 2.93E+06 4.56E+06
$\text{CoSpaDi}^{\dagger}$ x 0.5174 0.2635 0.0002 0.2538 0.2534 0.2100 0.2182 0.2413 0.2447 3.33E+05 2.16E+06
SVD-LLM v 0.6213 0.3643 0.2443 0.3603 0.2509 0.6490 0.2900 0.2298 0.3762 1.69E+02 1.70E+02
CoSpaDi v 0.2 0.6605 0.4288 0.3839 0.3994 0.2602 0.7160 0.3167 0.2483 0.4267 6.37E+01 3.51E+01
SVD x 0.5234 0.2564 0.0002 0.2416 0.2346 0.1950 0.2191 0.2698 0.2425 1.09E+06 3.92E+06
$\text{CoSpaDi}^{\dagger}$ x 0.5049 0.2626 0.0002 0.2449 0.2611 0.2180 0.2153 0.2552 0.2453 2.06E+05 4.34E+06
SVD-LLM v 0.5565 0.3006 0.0910 0.3047 0.2150 0.4590 0.2583 0.2323 0.3022 5.90E+02 2.47E+03
CoSpaDi v 0.3 0.5691 0.3241 0.1822 0.3194 0.2210 0.5670 0.2804 0.2308 0.3368 2.89E+02 6.59E+02
SVD x 0.5277 0.2590 0.0001 0.2386 0.2133 0.1990 0.2220 0.2693 0.2411 1.23E+06 4.17E+06
$\text{CoSpaDi}^{\dagger}$ x 0.5103 0.2626 0.0001 0.2546 0.2688 0.2130 0.2115 0.2549 0.2470 3.05E+06 3.67E+07
SVD-LLM v 0.5180 0.2725 0.0126 0.2685 0.2287 0.3230 0.2440 0.2297 0.2621 1.58E+03 3.30E+04
CoSpaDi v 0.4 0.5348 0.2824 0.0380 0.2778 0.2295 0.3690 0.2402 0.2310 0.2753 7.96E+02 9.23E+03
SVD x 0.5392 0.2566 0.0001 0.2441 0.2039 0.2040 0.2077 0.2690 0.2406 1.22E+06 1.88E+07
$\text{CoSpaDi}^{\dagger}$ x 0.5120 0.2607 0.0001 0.2572 0.2654 0.2280 0.2211 0.2704 0.2519 6.09E+05 1.08E+07
SVD-LLM v 0.5109 0.2663 0.0004 0.2609 0.2594 0.2610 0.2392 0.2303 0.2536 3.13E+03 1.03E+05
CoSpaDi v 0.5 0.5169 0.2703 0.0031 0.2626 0.2398 0.2950 0.2421 0.2326 0.2578 1.80E+03 7.26E+04
Quantization of Sparse Coefficient Matrix We investigated the quantization of the coefficient matrix $\mathbf{S}$ in the post-training regime to leverage the memory add-on due to the requirement of storing the indices of the nonzero values. We utilized a naive mantissa bit truncation of the original bfloat16 coefficient values. In Table 1 we report how truncation affects the performance at different compression levels.
According to the obtained results, truncation of 2 bits leads to negligible drop. Thus, in all our experiments CoSpaDi consists of bfloat16 dictionary and coefficient matrices but is evaluated with the truncated version of the sparse coefficient matrix using 14-bits.
@@ 4.3 Main Results
Per-Layer Scenario:
In this section we apply CoSpaDi for each projection layer independently and compare it against current sota training-free low-rank method – SVD-LLM Wang et al. (2025a).
(a) Accuracy & Perplexity for LLaMA3.2-1B
(b) Accuracy & Perplexity for Qwen3-0.6B
Figure 3: Average benchmark accuracy and WikiText perplexity for (a) LLaMA-3.2-1B and (b) Qwen-3 0.6B using SVD-LLM and CoSpaDiwith respect to compression ratio.
From the provided plots CoSpaDi significantly outperforms SVD-LLM in both avg. accuracy and perplexity for small Llama3 and Qwen3 models. We further investigate how our method scales to larger models of the same families with 8B parameters. The related results are provided in Table 4.3.
We have also performed comparisons on the same models and compression levels against other training-free compression strategies. In particular we compare CoSpaDi against ASVD Yuan et al. (2023), ReplaceMe Shopkhoev et al. (2025) and LLMPruner Ma et al. (2023), with the last two being sota pruning methods. These results are provided in Appendix A.6
Table 3: Performance comparison of CoSpaDi vs sota SVD-LLM on Llama3-8B and Qwen3-8B at different compression levels on different benchmarks. Best results are highlighted with bold.
@@
SVD-LLM 0.7106 0.5837 0.5925 0.5551 0.3396 0.8640 0.3550 0.3262 0.5408 4.07E+01 1.09E+01
CoSpaDi 0.2 0.7519 0.6649 0.7380 0.6654 0.4155 0.8950 0.3818 0.4282 0.6176 1.97E+01 4.27E+00
SVD-LLM 0.6578 0.4640 0.3806 0.4188 0.2765 0.7000 0.3177 0.2715 0.4358 1.47E+02 6.09E+01
CoSpaDi 0.3 0.7051 0.5620 0.6128 0.5421 0.3353 0.8570 0.3617 0.3224 0.5373 4.46E+01 9.18E+00
SVD-LLM 0.6028 0.3450 0.1143 0.3237 0.2449 0.4420 0.2574 0.2305 0.3201 5.49E+02 1.30E+03
CoSpaDi 0.4 0.6371 0.4138 0.3029 0.3914 0.2662 0.6850 0.3053 0.2538 0.4069 1.84E+02 1.19E+02
SVD-LLM 0.5381 0.2790 0.0041 0.2731 0.2389 0.2660 0.2230 0.2297 0.2565 1.37E+03 2.80E+04
CoSpaDi 0.5 0.5642 0.3102 0.0390 0.2988 0.2218 0.3770 0.2287 0.2301 0.2837 8.10E+02 7.64E+03
@@
SVD-LLM 0.7383 0.6391 0.6218 0.6869 0.4573 0.9010 0.4048 0.5473 0.6246 2.06E+01 6.40E+00
CoSpaDi 0.2 0.7650 0.6804 0.6559 0.7218 0.4889 0.9320 0.4067 0.6075 0.6573 1.81E+01 4.88E+00
SVD-LLM 0.7035 0.5522 0.5377 0.5926 0.3712 0.8720 0.3837 0.4482 0.5576 2.74E+01 1.07E+01
CoSpaDi 0.3 0.7242 0.6050 0.6255 0.6385 0.4121 0.8840 0.3952 0.5127 0.5997 2.29E+01 6.29E+00
SVD-LLM 0.6627 0.4458 0.3790 0.4503 0.2807 0.7730 0.3531 0.2901 0.4544 4.30E+01 3.59E+01
CoSpaDi 0.4 0.6888 0.4898 0.4991 0.4941 0.2986 0.8200 0.3675 0.3660 0.5030 3.59E+01 1.47E+01
SVD-LLM 0.6023 0.3543 0.1956 0.3363 0.2278 0.6670 0.2909 0.2305 0.3631 8.72E+01 2.69E+02
CoSpaDi 0.5 0.6159 0.3798 0.2791 0.3565 0.2295 0.6860 0.3167 0.2305 0.3868 6.82E+01 1.10E+02
Cross-Layer Scenario
In this section, we validate CoSpaDi on the cross-layer compression scenario, where a common dictionary is shared across different layers. To ensure a fair comparison, we adopt the identical layer selection and grouping procedure introduced in Basis Sharing Wang et al. (2024) (for a detailed description we refer to Appendix A.7) and compare the performance of the two methods on the Llama2-7B model Touvron et al. (2023b). As shown in Table LABEL:tab:main_results_group, CoSpaDi consistently outperforms Basis Sharing by a large margin across all benchmarks and compression rates (0.2–0.5). This performance gap suggests that CoSpaDi’s sparsity-aware, dictionary-based decomposition better preserves task-critical features under aggressive compression.
We believe that a more sophisticated layer grouping strategy, based on feature map similarity or end-to-end differentiable search, could work better with CoSpaDi and further widen this performance margin, but we leave this as a future research direction. We provide details for $k$ (dictionary size) and $s$ (sparsity) parameters in the Appendix A.7.

### v3

@@ 4.2.1 Capacity Allocation: the $k/s$ Ratio
Figure 2: Dual-axis plot showing average accuracy (  solid lines, left axis) and WikiText perplexity (--- dashed lines, right axis, inverted logarithmic scale) as functions of $\rho$ for Llama3.2-1B under three compression levels: 0.2, 0.3 and 0.4.
Figure 3: Dual-axis plot showing relative average accuracy (  solid lines, left axis) and relative WikiText perplexity (--- dashed lines, right axis) as functions of $\mathbf{S}$ bitwidth for Llama3-8B under three compression levels: 0.2, 0.3 and 0.4.
4.2.2 Data-Free vs. Data-Aware Compression
A core motivation for dictionary learning is that its union-of-subspaces form can be more expressive than a single low-dimensional subspace regardless of whether calibration is used. We therefore compare: (i) data-free truncated SVD vs. CoSpaDi (dictionary learning in weight space, without whitening), and (ii) data-aware SVD-LLM vs. CoSpaDi (activation-aware via whitening). Table 2 summarizes results on LLaMA-3.2-1B across CR 0.2–0.4. In both regimes, the dictionary-based factorization is consistently stronger than the corresponding low-rank baseline, and activation-aware CoSpaDi yields the best overall trade-off.
@@
Llama3.2 1B 74.5 63.7 63.0 60.5 36.2 88.3 37.8 37.0 57.6 11.6E+00 5.7E+00
SVD
52.0 25.7 0.0 24.7 22.1 20.0 21.4 25.4 23.9 2.9E+06 4.6E+06
$\text{CoSpaDi}^{\dagger}$
51.7 26.4 0.0 25.4 25.3 21.0 21.8 24.1 24.5 3.3E+05 2.2E+06
SVD-LLM
62.1 36.4 24.4 36.0 25.1 64.9 29.0 23.0 37.6 1.7E+02 1.7E+02
CoSpaDi
0.2 66.1 42.9 38.4 39.9 26.0 71.6 31.7 24.8 42.7 6.4E+01 3.5E+01
SVD
52.3 25.6 0.0 24.2 23.5 19.5 21.9 27.0 24.3 1.1E+06 3.9E+06
$\text{CoSpaDi}^{\dagger}$
50.5 26.3 0.0 24.5 26.1 21.8 21.5 25.5 24.5 2.1E+05 4.3E+06
SVD-LLM
55.7 30.1 9.1 30.5 21.5 45.9 25.8 23.2 30.2 5.9E+02 2.5E+03
CoSpaDi
0.3 56.9 32.4 18.2 31.9 22.1 56.7 28.0 23.1 33.7 2.9E+02 6.6E+02
SVD
52.8 25.9 0.0 23.9 21.3 19.9 22.2 26.9 24.1 1.2E+06 4.2E+06
$\text{CoSpaDi}^{\dagger}$
51.0 26.3 0.0 25.5 26.9 21.3 21.2 25.5 24.7 3.1E+06 3.7E+07
SVD-LLM
51.8 27.3 1.3 26.9 22.9 32.3 24.4 23.0 26.2 1.6E+03 3.3E+04
CoSpaDi
0.4 53.5 28.2 3.8 27.8 23.0 36.9 24.0 23.1 27.5 8.0E+02 9.2E+03
4.2.3 Dictionary Learning Solver Variants: Exact K-SVD vs. Power-K-SVD vs. MOD
The main bottleneck in the alternating minimization process is the dictionary update using K-SVD as it updates a single atom at a time. To reduce its computational cost we investigate the usage of truncated PCA and power iterations for solving rank-1 problems as well as the well-known Method of Optimal Directions (MOD) which updates in a single step the entire dictionary instead of individual atoms sequentially.
Table 2 reports both quality (Avg. accuracy, WikiText perplexity) and wall-clock compression time at a representative 0.2 CR on LLaMA-3.2-1B.Wall-clock is measured for compressing all targeted projection matrices with identical hardware, batching, and stopping criteria. This comparison isolates the practical speed–quality trade-off of the solver choice while keeping the factorization form $(k,s)$ fixed. For our further experiments we utilized K-SVD using power-iterations as it achieves good performance at a reasonable wall-clock time for compression. We provide a more thorough ablation study on the number of power iterations and of the alternating steps in Appendix A.7
Table 2: Solver comparison for dictionary update on LLaMA-3.2-1B at 0.2 compression ratio (CR) and fixed $\rho=2$, 60 alternating minimization iterations. We report accuracies, perplexity as well as wall-clock compression time in minutes using a single A100. Best results are highlighted with bold.
@@
Method time PIQA Hella Swag LAMBADA ARC-e ARC-c SciQ Race MMLU Avg. Wiki Text LAMBADA
Llama3.2 1B 74.5 63.7 63.0 60.5 36.2 88.3 37.8 37.0 57.6 11.6E+00 5.7E+00
MOD 222.2 62.9 39.7 30.4 36.3 25.7 67.6 27.9 23.0 39.2 1.1E+02 8.7E+01
K-SVD (PCA) 902.1 64.4 39.9 33.7 40.5 25.9 72.3 31.6 24.4 41.6 9.6E+01 6.2E+01
K-SVD (power) 646.8 75.2 66.5 73.8 66.5 41.6 89.5 38.2 42.8 61.8 2.0E+01 4.3E+00
4.2.4 Quantizing Coefficients and Removing the Mask Term
Sparse factorization introduces metadata overhead for storing nonzero locations. Our default implementation uses a packed binary mask (one bit per entry) plus bf16 values for the $sd_{2}$ nonzeros. We additionally study post-training coefficient quantization by truncating bf16 mantissa bits. Plots on Figure 3 shows that truncating 2 mantissa bits yields negligible degradation across tested CRs.
Beyond robustness, coefficient quantization enables a simplified storage accounting in which locations can be treated as fixed by construction (e.g., using a deterministic pattern per column or a reproducible pseudo-random seed), removing the explicit mask term from the compression ratio. Under this view, the effective storage becomes
$\hat{\gamma}^{\mathrm{SD}}=1-\frac{d_{1}k+sd_{2}}{d_{1}d_{2}}.$ (14)
Unless stated otherwise, we evaluate CoSpaDi with bf16 dictionaries and truncated (14-bit) coefficients, thus providing CR according to Equation 14.
@@ 4.3 Main Results
4.3.1 Per-Layer Compression
We first compress each projection matrix independently and compare CoSpaDi to SVD-LLM over a range of CRs. Figure 4 summarizes the accuracy–compression and perplexity–compression curves for small models (LLaMA-3.2-1B and Qwen-3-0.6B). For larger models, Table 3 reports results on LLaMA-3-8B, Qwen-3-8B and Qwen-3-14B at CR 0.2–0.4. Across model families and budgets, CoSpaDi consistently improves both average benchmark accuracy and perplexity at matched CR, indicating that the union-of-subspaces factorization better preserves task-relevant directions than a single low-rank basis. We provide additional benchmarks in the Appendix A.6.
Figure 4: Average benchmark accuracy and WikiText perplexity for LLaMA-3.2-1B and Qwen-3 0.6B using SVD-LLM and CoSpaDi with respect to compression ratio.
Table 3: Performance comparison of CoSpaDi vs sota SVD-LLM on Llama3-8B, Qwen3-8B and Qwen3-14B at different compression levels on different benchmarks. Best results are highlighted with bold.
@@ 4.3.1 Per-Layer Compression
Llama3 8B 80.7 79.1 75.6 77.7 53.5 93.9 40.3 62.2 70.4 7.3E+00 3.1E+00
SVD-LLM 71.1 58.4 59.3 55.5 34.0 86.4 35.5 32.6 54.1 4.1E+01 1.1E+01
CoSpaDi 0.2 75.2 66.5 73.8 66.5 41.6 89.5 38.2 42.8 61.8 2.0E+01 4.3E+00
SVD-LLM 65.8 46.4 38.1 41.9 27.7 70.0 31.8 27.2 43.6 1.5E+02 6.1E+01
CoSpaDi 0.3 70.5 56.2 61.3 54.2 33.5 85.7 36.2 32.2 53.7 4.5E+01 9.2E+00
SVD-LLM 60.3 34.5 11.4 32.4 24.5 44.2 25.7 23.1 32.0 5.5E+02 1.3E+03
CoSpaDi 0.4 63.7 41.4 30.3 39.1 26.6 68.5 30.5 25.4 40.7 1.8E+02 1.2E+02
Qwen3 8B 77.7 74.9 64.1 80.7 56.7 95.7 40.9 73.0 70.5 1.2E+01 4.6E+00
SVD-LLM 73.8 63.9 62.2 68.7 45.7 90.1 40.5 54.7 62.5 2.1E+01 6.4E+00
CoSpaDi 0.2 76.5 68.0 65.6 72.2 48.9 93.2 40.7 60.8 65.7 1.8E+01 4.9E+00
SVD-LLM 70.4 55.2 53.8 59.3 37.1 87.2 38.4 44.8 55.8 2.7E+01 1.1E+01
CoSpaDi 0.3 72.4 60.5 62.6 63.9 41.2 88.4 39.5 51.3 59.97 2.3E+01 6.3E+00
SVD-LLM 66.3 44.6 37.9 45.0 28.1 77.3 35.3 29.1 45.4 4.3E+01 3.6E+01
CoSpaDi 0.4 68.9 49.0 49.9 49.4 29.9 82.0 36.8 36.6 50.3 3.6E+01 1.5E+01
Qwen3 14B 79.9 78.9 67.9 82.8 60.2 96.5 43.3 77.2 73.3 1.1E+01 3.7E+00
SVD-LLM 76.2 67.6 69.1 69.8 46.8 91.0 42.9 62.5 65.8 1.8E+01 4.1E+00
CoSpaDi 0.2 77.3 71.7 72.0 73.3 51.3 91.9 43.1 65.8 68.3 1.6E+01 3.4E+00
SVD-LLM 72.0 59.7 61.5 62.6 39.3 88.9 41.5 52.0 59.7 2.3E+01 6.5E+00
CoSpaDi 0.3 74.6 64.2 69.0 65.3 43.6 90.7 42.5 58.0 63.5 2.0E+01 4.2E+00
SVD-LLM 67.4 48.7 46.0 48.5 30.7 80.6 36.7 36.8 49.4 3.5E+01 1.8E+01
CoSpaDi 0.4 69.5 52.3 57.3 53.2 33.0 84.4 37.5 43.6 53.9 3.3E+01 8.7E+00
4.3.2 Comparison with Pruning
We additionally compare CoSpaDi to recent structured pruning approaches on the same models and compression ratios. Specifically, we evaluate against ReplaceMe (Shopkhoev et al., 2025) and LLM-Pruner (Ma et al., 2023), which are representative state-of-the-art methods that reduce parameters by removing or replacing structured components (e.g., blocks/layers) rather than individual weights. Results are reported in Table 4.
We do not include unstructured pruning baselines in this comparison because their practical memory savings depend on sparse storage formats and hardware support; at moderate sparsity levels (e.g., $\approx$50%), the overhead of storing indices and irregular access patterns often diminishes or eliminates effective compression and speedups in standard dense inference pipelines.
Table 4: Comparison of the proposed CoSpaDi method with state-of-the-art structured pruning methods ReplaceMe Shopkhoev et al. (2025) and LLM-Pruner Ma et al. (2023) on Llama3 8B under different compression ratios. We report accuracy on different benchmarks as well as its average and perplexity. Best results are highlighted with bold
@@ 4.3.2 Comparison with Pruning
Llama3 8B 80.7 79.1 75.6 77.7 53.5 93.9 40.3 62.2 70.4 7.3E+00 3.1E+00
ReplaceMe 0.22 73.1 65.7 42.1 65.9 43.7 86.4 35.4 51.7 58.0 3.4E+01 2.0E+01
LLM-Pruner 75.5 67.5 51.0 62.1 36.6 87.8 35.1 25.0 55.1 1.6E+01 1.1E+01
CoSpaDi 0.2 75.2 66.5 73.8 66.5 41.6 89.5 38.2 42.8 61.8 2.0E+01 4.3E+00
ReplaceMe 0.31 66.6 53.8 24.0 50.7 37.9 77.3 34.0 30.6 46.9 6.7E+01 1.3E+02
LLM-Pruner 67.3 45.1 20.9 45.4 28.8 63.4 30.1 22.9 40.5 3.8E+01 2.2E+02
CoSpaDi 0.3 70.5 56.2 61.3 54.2 33.5 85.7 36.2 32.2 53.7 4.5E+01 9.2E+00
ReplaceMe 0.41 61.7 44.3 9.8 37.4 27.5 60.4 31.6 26.4 37.4 2.3E+02 1.8E+03
LLM-Pruner 50.3 25.8 1.5 26.4 25.8 28.1 21.8 23.2 25.4 $\infty$ 5.7E+05
CoSpaDi 0.4 63.7 41.4 30.3 39.1 26.6 68.5 30.5 25.4 40.7 1.8E+02 1.2E+02
4.3.3 Grouped (Cross-Layer) Dictionary Sharing
Next, we evaluate grouped compression where a single dictionary is shared across multiple layers of the same projection type. For a fair and controlled comparison, we adopt the exact layer grouping and evaluation protocol of Basis Sharing, i.e., we share parameters across the same layer sets and report results under the same compression budgets. Table 5 reports results on LLaMA-2-7B across CR 0.2–0.4. We additionally included per-layer counterparts to show the benefits of grouping strategy and exploiting inter-layer redundancies. At all budgets, CoSpaDi substantially outperforms Basis Sharing as well as per-layer baselines, suggesting that cross-layer sharing is most effective when paired with sparse codes that allow each column (and each layer) to select its own subset of atoms. We note that CoSpaDi does not rely on this particular grouping and could benefit from more adaptive sharing strategies (e.g., data-driven grouping or partially shared dictionaries), which we leave for future work.
Table 5: Performance comparison of CoSpaDi vs Basis Sharing (Wang et al., 2025a) and per-layer counterparts on Llama2-7B under different compression levels on various benchmarks. Best results are highlighted with bold.
@@ 4.3.3 Grouped (Cross-Layer) Dictionary Sharing
Llama2 7B 78.0 57.1 73.7 76.1 43.3 93.9 39.2 40.9 62.8 8.7 3.4
SVD-LLM 73.3 60.4 64.0 55.1 32.3 79.7 36.6 25.5 53.4 20.2 6.3
CoSpaDi (per-layer) 74.3 64.8 70.2 61.6 35.3 85.8 39.0 28.2 57.4 15.8 4.5
Basis Sharing 70.1 43.4 63.0 66.5 33.7 91.0 34.9 24.8 53.4 15.2 7.2
CoSpaDi (grouped) 0.2 74.6 66.3 71.3 67.4 39.3 88.8 38.3 27.2 59.2 11.7 4.3
SVD-LLM 68.4 52.2 51.5 46.8 27.8 75.9 34.7 23.5 47.6 34.0 13.5
CoSpaDi (per-layer) 71.2 57.4 62.7 53.5 30.6 81.6 36.8 26.9 52.6 23.6 6.9
Basis Sharing 65.6 37.6 53.5 58.8 27.3 86.8 32.6 23.2 48.2 22.2 13.6
CoSpaDi (grouped) 0.3 70.8 58.5 64.5 61.9 34.3 87.6 35.9 24.0 54.7 15.4 6.3
SVD-LLM 63.2 42.7 33.1 37.7 24.6 70.2 32.2 23.0 40.8 86.5 67.1
CoSpaDi (per-layer) 66.2 48.4 47.2 44.3 26.2 76.7 33.4 23.4 45.7 49.2 20.9
Basis Sharing 60.6 33.1 41.1 49.5 23.6 83.0 29.7 23.1 42.95 39.5 37.7
CoSpaDi (grouped) 0.4 64.3 48.0 51.8 52.3 29.9 81.5 33.2 23.0 48.0 25.2 14.6
5 Open Issues and Future Work
Solver efficiency and scalability. CoSpaDi uses alternating minimization with OMP for sparse coding and K-SVD-style sequential atom updates. This is simple and effective but can be expensive due to repeated sparse coding and per-atom updates, especially for large matrices and aggressive compression. Faster variants (e.g., more parallel updates, better batching, or alternative sparse coding) are promising; our ablations already show that approximate rank-1 updates (e.g., power iterations) offer a favorable accuracy–time trade-off.
Hyperparameter selection and dynamic budget allocation. We parameterize CoSpaDi via $(k,s)$ and $\rho=k/s$, but fix $\rho$ and use uniform budgets for simplicity. Prior low-rank work shows that layer-/type-wise sensitivity can improve quality at a fixed global budget. Adapting dynamic allocation to dictionary learning–e.g., choosing $(k_{\ell},s_{\ell})$ per layer, projection type, or group based on functional sensitivity–is a promising direction.
Grouped sharing design choices. Grouped performance depends on how layers are clustered and which projections share a dictionary. We follow a simple grouping aligned with prior cross-layer sharing, while more adaptive grouping based on activation/weight similarity could yield additional gains.
Practical speedups depend on kernel support. CoSpaDi yields structured sparsity exploitable by sparse–dense computation, but end-to-end speedups depend on sparse kernel efficiency, indexing overhead, and reuse of intermediate products. Encouragingly, recent work reports strong acceleration for dense–sparse matmul even at modest sparsity, narrowing the gap between parameter reduction and wall-clock benefits (Macko & Boža, 2025). Realizing consistent speedups will require integration into optimized inference runtimes and hardware-aware kernels for the resulting computation patterns.
Beyond the current instantiation. Our goal is to introduce activation-aware sparse dictionary learning as a new post-training compression paradigm beyond single-subspace (low-rank) approximations. Accordingly, we use a direct optimization pipeline. Many advances from low-rank compression–richer calibration objectives, better sensitivity measures, and refined allocation strategies–could be transferred to this setting without changing the core factorization.
6 Conclusion
We presented CoSpaDi, a training-free LLM compression framework based on sparse dictionary learning. CoSpaDi factorizes projection matrices into a dense dictionary and column-sparse codes, moving from a single shared low-dimensional subspace to a union-of-subspaces representation. Using a small calibration set, an activation-derived Gram orthonormalization converts functional output reconstruction into a tractable dictionary learning objective solved by alternating sparse coding and dictionary updates.
Across multiple Llama and Qwen models, CoSpaDi improves the accuracy–compression and perplexity–compression trade-offs over strong SVD-based baselines in both per-layer and grouped (shared-dictionary) settings at 20–40% compression ratios, and composes naturally with post-training coefficient quantization. More broadly, our results highlight dictionary learning as a flexible paradigm for post-training compression and suggest that many advances from low-rank methods (e.g., dynamic allocation and improved calibration objectives) can be transferred to this richer factorization family.